OpenStudy (anonymous):

Find the y intercept and the x intercept (ordered pairs) -10 x + y = 6

jimthompson5910 (jim_thompson5910):

x intercept -10 x + y = 6 -10x + 0 = 6 -10x = 6 x = -6/10 x = -3/5 So the x intercept is (-3/5, 0) y intercept -10x + y = 6 -10(0) + y = 6 0 + y = 6 y = 6 So the y intercept is (0,6)